# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v2

@@ E.2 Pixel-wise Classification Baselines
  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model consisted of 200 trees, with other parameters set to standard values.
@@ E.3 Patch-wise Semantic Segmentation
  • For other foundation models that are inherently patch-based (e.g., Prithvi [141], Satlas [18]), their encoders already process image patches. Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output. We did not freeze the encoders during the finetuning.
@@ E.4 Embedding Space Analysis
The 2D visualizations shown in the main text (Figure 2 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [103] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embeddings map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for TESSERA, PRESTO, and GSE embeddings.
@@
    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighboring cluster. The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
@@
    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$. Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where clusters are compact and well separated.
Figure 11: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labeled ’Embedding’) and the final predicted segmentation map (labeled ’Predicted’).
@@ F.1 Study Areas and Ground Truth Datasets
Canopy height prediction was evaluated in a 5$\times$6 km area in the Danum Valley, Borneo. This region is covered in tropical rainforest dominated by the fast-growing dipterocarp family, which hosts some of the tallest trees in the tropics [135, 120]. Ground truth data was derived from airborne laser scanning (ALS) performed in February 2020 [33]. ALS data provides detailed 3D point clouds of vegetation and terrain structure, which were processed using the  lspikefree algorithm [51] to estimate top-of-canopy height at 1 m resolution. These canopy height rasters were then downsampled using mean pooling to 10 m to match the resolution of the TESSERA representations.
F.2 Modeling Methods
@@ F.3 Baselines
To benchmark performance against traditional remote sensing approaches, we compared the predictions against three global canopy height products: Potapov et al. [2021][121] (GLAD), Lang et al. [2023][84] (ETH), and Tolan et al. [2024][148] (Meta) as well as a regional map produced by Lang et al. [2021][85]. These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region.
Finally, we trained the U-Net model using Sentinel-2 input as a baseline. This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery. Cloud masking and compositing were performed in GEE, resulting in 12 channel input images (band XYZ excluded), with one channel for each Sentinel-2 spectral band. Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
@@ G.1 Study Areas and Ground Truth Datasets
For fire disturbance detection, two 15 × 15 km tiles in California were selected at random from the pool of training samples in the burn scars benchmarking dataset [76], referred to as Burn Area A and Burn Area B. Burn Area A comprises a heterogeneous landscape of forest, shrubland, agricultural land, and built infrastructure. In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha). Burn Area B experienced a single large fire ’Glass’ ($\sim$27,500 ha) in September 2020 that primarily affected forested land. In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland. To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [2].
The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and burn severity. Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset. Burn severity–defined as the measure of ecological damage caused by fire [55]–was derived from categorical and continuous severity metrics provided by MTBS. To obtain burn severity categories of low, medium and high, as well as continuous burn severity, we used MTBS burn severity annualized dataset as well as individual severity maps for each fire. MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images. Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [55]. Since MTBS datasets are derived from Landsat, they are provided at 30-meter resolution, and were upsampled to 10-meters to match the resolution of TESSERA representations. The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [57].
@@ G.2 Dimensionality Reduction
We used UMAP (Uniform Manifold Approximation and Projection) [103] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires. We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections. To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs. unburned), land cover class, or burn severity. These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest. Each Burn Area contains over two million 10-meter pixels. To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction. UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space. UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
@@ H.1 Dataset
For the AGB regression task, we used the ground truth data from the Biomassters dataset [111]. Patch coordinates and collection years [111] were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [56].
For the SpectralGPT baseline, we used the preprocessed satellite data from the Biomassters dataset as input, since the codebase for SpectralGPT [67] did not include a preprocessing pipeline at the time of our analysis.
@@ H.2 UNet model architecture
To obtain AGB estimates, we trained a convolutional neural network based on a UNet architecture [128], using TESSERA or GSE embeddings  [56] as input representations. The model received 128$\times$256$\times$256 and 64$\times$256$\times$256 input patches for TESSERA and GSE, respectively, where the first dimension denotes the number of channels.
The network followed a standard UNet structure with two encoder and decoder stages. The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling. A 1024-channel bottleneck block was used prior to upsampling. The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double convolution blocks. A final 1×1 convolution produced the single-channel regression output. The model had approximately 30 million trainable parameters. The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
Training was performed with the Adam optimizer (learning rate = 1e-4, batch size = 4). Finetuning on the full dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours.
@@ H.3 Baselines
In addition to GSE, we compared TESSERA’s performance to two other baselines: the Biomassters winner [111] and the SpectralGPT foundation model [67] combined with an UPerNet [179].
For SpectralGPT, we used the implementation provided in Ref. [101], with the Biomassters satellite data as input and an UPerNet as the regression head. This setup was found in Ref. [101] to outperform other tested foundation models. To ensure consistency in the limited-label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation. Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [101] as it resulted in slightly better performance. Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
To evaluate the Biomassters winning model, we downloaded the pre-trained model weights and ran inference on the test set. Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model. According to the Biomassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days – highlighting the greater time efficiency of TESSERA, which required only five hours.
@@ Appendix I Downstream Task: Stocking indices in voluntary carbon markets
I.1 Relative height modeling
Participants in the voluntary carbon market can use a remotely sensed ”stocking index” to quantify project additionality. The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
@@ I.1 Relative height modeling
To model a suitable stocking index from the 128-dimensional TESSERA embeddings, 110,000 relative heights RH95 and RH10 from 2022 were extracted from the Sentinel-2 tile 22MCT in Para Sate, Brazil. From the 10 m pixel falling at the center of the full-waveform footprint, the 128 TESSERA embeddings were extracted as predictors. After data extraction, the relative height differences (dRH = RH95-RH10) were used as the target variable and the 128 TESSERA embeddings as predictors. A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index. To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training. The TerraClass Amazonia 2022 land cover map [146] was used to distinguish between forest and non-forest.
For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC). Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
@@ I.2 Stocking index baselines
For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [156], ESA-CCI AGB (100 m) [132], Meta CHM (1 m)  [148] and GTDX AGB (25 m) [42]. Using the same shapefiles delineating the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
@@
The core of the tool is the interactive labeling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labeled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in our github repository [152].
@@
Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10m resolution (10m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: $\checkmark$= Yes, $\times$= No, ? = Unknown, p= Potentially true, .. = N/A.
@@
Galileo (2025) [150] p p $\checkmark$ ? $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v3

@@ E.2 Pixel-wise Classification Baselines
  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
@@ E.3 Patch-wise Semantic Segmentation
  • For other foundation models that are inherently patch-based (e.g., Prithvi [143], Satlas [19]), their encoders already process image patches. Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output. We did not freeze the encoders during the fine-tuning.
@@ E.4 Embedding Space Analysis
The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [102] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
@@
    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster. The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
@@
    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$. Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
@@ F.1 Study Areas and Ground Truth Datasets
Canopy height prediction was evaluated in a 5$\times$6 km area in the Danum Valley, Borneo. This region is covered in tropical rainforest dominated by species in the Dipterocarpaceae family, which hosts some of the tallest trees in the tropics [137, 119]. Ground truth data was derived from airborne laser scanning (ALS) performed in February 2020 [34]. ALS data provides detailed 3D point clouds of vegetation and terrain structure, which were processed using the  lspikefree algorithm [52] to estimate top-of-canopy height at 1 m resolution. These canopy height rasters were then downsampled using mean pooling to 10 m to match the resolution of the TESSERA representations.
F.2 Modelling methods
@@ F.3 Baselines
To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al. [2021][120] (GLAD), Lang et al. [2023][84] (ETH), and Tolan et al. [2024][150] (Meta) as well as a regional map produced by Lang et al. [2021][85]. These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
Finally, we trained the U-Net model using Sentinel-2 input as a baseline. This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery. Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band. Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
@@ G.1 Study Areas and Ground Truth Datasets
For fire disturbance detection, two 15$\times$15 km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [77], referred to as Burn Area A and Burn Area B. Each Burn Area contains over two million 1 m pixels. Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure. In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha). Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land. In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland. To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [2].
The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and severity of the burn. Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset. Burn severity, defined as the measure of ecological damage caused by fire [56], was derived from categorical and continuous severity metrics provided by MTBS. To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire. MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images. Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [56]. Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations. The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [58]. For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high). The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
@@ G.2 Dimensionality Reduction
We used UMAP (Uniform Manifold Approximation and Projection) [102] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires. We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections. To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs. unburned), land cover class, or burn severity. These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest. To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction. UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space. UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
G.3 Label Efficiency Analysis and Baselines
We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available. The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set. We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%. The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance. Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
We compared three types of input embeddings for the classification task. The first was TESSERA, using embeddings for 2020. The second was a task-specific composite baseline, constructed using prior knowledge of the fire event. This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector. The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [57]. GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
@@ H.1 Dataset
For the AGB regression task, we used the ground truth data from the BioMassters dataset [110]. This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a 2,560$\times$2,560  m area at 10 meter pixel resolution. Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre. Patch coordinates and collection years–provided by the authors of Ref. [110] upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [57].
@@ H.2 UNet model architecture
To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [129], using TESSERA or GSE embeddings  [57] as input. The model received 256$\times$256$\times$128 and 256$\times$256$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
The network followed a standard UNet structure with two encoder and decoder stages. The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling. A 1024-channel bottleneck block was used prior to upsampling. The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks. A final 1×1 convolution produced the single-channel regression output. The model had approximately 30 million trainable parameters. The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
Following the evaluation protocol in Ref. [110], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
@@ H.3 Baselines
In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [110] and the SpectralGPT foundation model [68] combined with a UPerNet [181].
For SpectralGPT, we used the implementation provided in Ref. [100], with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE [133] for temporal aggregation, and a UPerNet as the regression head. This setup was found in [100] to outperform other tested foundation models on the BioMassters dataset. To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation. Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [100] because it resulted in slightly better performance. Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range. Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model. According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
@@ Appendix I Downstream Task: Stocking indices in Voluntary Carbon Markets
I.1 Relative height modelling
Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events. The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
@@ I.1 Relative height modelling
To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [42] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil. From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables. A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index. The TerraClass Amazonia 2022 land cover map [148] was used to distinguish between forest and non-forest. To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
@@ I.2 Stocking index baselines
For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC). Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA. For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [158], ESA-CCI AGB (100 m) [134], Meta CHM (1 m)  [150], and GTDX AGB (25 m) [43]. Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
@@
The core of the tool is the interactive labelling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labelled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in our github repository [154].
@@
Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: $\checkmark$= Yes, $\times$= No, ? = Unknown, p= Potentially true, .. = N/A.
@@
Galileo (2025) [152] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$